# Large Language Models Are Unconscious of Unreasonability in Math Problems

### v1

@@ Abstract
Large language models (LLMs) demonstrate substantial capabilities in solving math problems. However, they tend to produce hallucinations when given questions containing unreasonable errors. In this paper, we study the behavior of LLMs when faced with unreasonable math problems and further explore their potential to address these problems. First, we construct the Unreasonable Math Problem (UMP) benchmark to examine the error detection ability of LLMs. Experiments show that LLMs are able to detect unreasonable errors, but still fail in generating non-hallucinatory content. In order to improve their ability of error detection and correction, we further design a strategic prompt template called Critical Calculation and Conclusion(CCC). With CCC, LLMs can better self-evaluate and detect unreasonable errors in math questions, making them more reliable and safe in practical application scenarios.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (ope, 2023; gem, 2023) have significantly advanced the progress of artificial intelligence. These models demonstrate impressive performance across a wide range of downstream tasks, including solving math problems like GSM8K (Cobbe et al., 2021) and MATH(Hendrycks et al., 2021).
Figure 1: An overview of our self-detection process. We urge the model to scrutinize the reasonability of each question, then directly answering reasonable ones while correcting and reevaluating unreasonable ones.
In scenarios such as intelligent education, an LLM with strong math ability can be a powerful and versatile tool.
@@
Figure 2: Example inputs and outputs with a) direct mathematical problem prompts and b) prompts that signal the potential unreasonableness of the math problems. When directly queried, the model methodically works through the problem to completion, sometimes producing illogical results, such as "12.5 trees". However, if the prompt includes hints about the problem’s reasonableness without any parameter modification, the model can effectively identify and highlight the nonsensical aspects of the query.
@@
Considering the strong error detection capability in LLMs, we further design a prompt template called Critical Calculation and Conclusion to stimulate and leverage the self-evaluation and critical thinking abilities in LLMs. To be specific, for each question, We urge the model to scrutinize the reasonability of each question. Then, it will directly answer reasonable questions, but correct and reevaluate unreasonable ones. An overview of this process is illustrated in Figure 1. With CCC, LLMs enhance their ability to identify and highlight unreasonable part within a problem. Simultaneously, this approach ensures the model’s performance remains unaffected when dealing with problems that are deemed reasonable.
Our contributions are summarized as follows: (1) We find that many LLMs’ performance significantly diminishes when encountering unreasonable problems, posing potential security risks. (2) We develop the Unreasonable Math Problem (UMP) benchmark to systematically assess the model’s ability in addressing unreasonable problems. (3) We introduce the CCC prompt template designed to stimulate the model’s self-detection capabilities. This method aids the model in identifying and rectifying unreasonable problems efficiently.
2 Related Work
2.1 LLMs as Math Problem Solver
2.2 Improve models’ inference ability
Methodologically, a variety of strategies have been developed to augment the model’s reasoning capabilities. For instance, the "chain of thought" approach, introduced by Wei et al. (2022), enables models to generate more accurate solutions by sequentially processing information and emulating human cognitive processes. Complex COT(Fu et al., 2023) prompts models with the most intricate problems and solutions, thereby encouraging the generation of sophisticated responses. Plan-and-Solve (Wang et al., 2023) method enhance model performance by first generating a plan or sequence of steps, which is then executed to arrive at a solution. Yu et al.’s (2023a) output-supervised verifiers refine the model’s capabilities by scrutinizing the generated output and emphasizing the most critical steps within the problem-solving process.
@@ 3 Explore how LLMs respond when faced with unreasonable problems
Overview Although the existing technology can achieve good results in solving mathematical reasoning problems, we assume that such techniques often rely on the model’s ability to internalize and apply procedural templates to novel problems, but not inherently prioritizes procedural execution over semantic understanding. As illustrated in figure 2, when confronted with nonsensical or unreasonable questions, the models are prone to default to the application of memorized formulas and procedures, thereby generationg some inaccurate and confusing responses. To thoroughly evaluate the performance of large language models in the face of the above phenomenon, we construct a benchmark named Unreasonable Math Problems, containing unreasonable questions and test it on different models.
@@ 3.1 The UMP Benchmark
Our benchmark comprises 100 mathematically unreasonable problems and 100 original reasonable problems. For Each problem has been systematically derived from a corresponding valid question sourced from the established GSM8K dataset (Cobbe et al., 2021), ensuring that the foundational structure of the problems remains stable. This subsection details the benchmark construction process.
@@ 3.1.1 Data Category
We identify and categorize five prevalent types of unreasonable math problems that commonly challenge large language models. These categories are: Undefined Variables, Illogical Scenarios, Incorrect Assumptions, Misinterpretation of Units, and Inconsistent Conditions. Below, we provide detailed interpretations for each of these categories to facilitate a deeper understanding of the issues that may arise when models are confronted with such problems.
@@
3.1.2 Data Generation
We adopt a methodology akin to that of MetaMath(Yu et al., 2023b), utilizing ChatGPT to facilitate the generation of data. By selecting a subset of problems from the GSM8K(Cobbe et al., 2021) dataset, we prompt the model to rephrase these questions. Through the use of carefully crafted prompts, we guide ChatGPT to intentionally alter the original problems, thereby producing the specific types of unreasonable questions required for our analysis. 111the prompt of generating unreasonable problems can be found in Appendix A
Each of our data pieces is made up of five parts. An original question $q$ paired with its corresponding answer $a$, along with a designated type of unreasonableness $t$ we aim to investigate, we systematically construct a rephrased, unreasonable variant $q^{{}^{\prime}}$ and provide an accompanying explanation $e$. Consequently, our benchmark comprises quintuples of the form
$(q,a,q^{{}^{\prime}},t,e)$
Example:Original question and Unreasonable question original question:Jenna and her mother picked some apples from their apple farm. Jenna picked half as many apples as her mom. If her mom got 20 apples, how many apples did they both pick? answer: 30 type: Incorrect Assumptions new question: Jenna and her mother picked some apples from their apple farm. Jenna picked half as many apples as her mom. If her mom got 15 apples, how many apples did they both pick? explanation:This question is unreasonable because picking half an apple is not practical in real-life scenarios (as apples are typically picked whole) In this case, if we adhere strictly to the question, they picked a total of 15 (mom’s apples) + 7.5 (Jenna’s apples) = 22.5 apples. The only unreasonable aspect might be the fractional apple, as it suggests that an apple was divided, which is not typical when picking apples from a tree.
3.1.3 Data Checking
To better test the model’s detection capabilities, we aim to make the unreasonable problems in the benchmark less conspicuous. Initial model-generated questions, particularly within the ’Incorrect Assumptions’ and ’Inconsistent Conditions’ categories, were often too evident in their unreasonableness, as can be seen in Figure 3. For instance, some questions contained overtly unreasonable statements within the problem statement itself, which is not what we want to achieve. Therefore, following the generation of questions by the LLM using our prompt, we conducted thorough manual reviews and edited the content. The alterations included changing variables and rephrasing, where necessary, to render the questions superficially reasonable while preserving their underlying unreasonable part.
Figure 3: This figure shows how we edit unreasonable questions for subtlety. On the left, we present the initial reasonable problem. The middle has the first draft by the LLM, which is clearly nonsensical with errors like fractional people. Then we manually edit these questions, resulting in the example on the right. which looks normal but has the number of men and women exceeds the total population, keeping the question secretly unreasonable.
3.2 Performance on different LLMs
3.2.1 Explore LLMs’ ability to judge whether the statements are reasonable
To assess the model’s response to unreasonable problems, it’s imperative to first verify its proficiency in recognizing unreasonable expressions. Without the capability to pinpoint inaccuracies directly, the model’s effectiveness in detecting embedded errors within questions becomes questionable. Our preliminary analysis reveals that errors such as Incorrect Assumptions and Misinterpretation of Units predominantly stemmed from common sense misunderstandings, e.g., the impossibility of negative values for human attributes or the recognition of a gram as a measure of weight. To further investigate, we devise a small benchmark targeting these categories, comprising 30 items each, to evaluate the model’s discernment of such irrational statements upon direct confrontation. The methodology employed is detailed in the Appendix C. The result in table 1 indicate that models, even those of a 6B size, generally possess the competency to assess the rationality of statements.
Model/Accuracy IA MU
ChatGLM3-6B 80 76.6
Llama2-7B-chat 80 73.3
Baichuan2-13B-chat 83.3 83.3
Qwen-Max 100 96.6
GPT-3.5 93.3 100
GPT-4 100 96.6
Table 1: The accuracy with which various models successfully identified irrationality when directly confronted with errors categorized as Incorrect Assumptions (IA) and Misinterpretation of Units (MU). For each category, we conduct tests on a benchmark comprising thirty instances. The results show that the model basically has enough commonsense knowledge to determine whether the statements are unreasonable.
3.2.2 Performance on UMP benchmark
We adopt the common Chain of Thought (CoT) method as a prompt, structured as follows: "As an AI assistant, you’re equipped to assist with mathematical problems. Q: [X]", where X denotes the specific mathematical problem in question. To evaluate the model’s output, we first look for keywords or expressions like "unreasonable". Next, we assess whether the model has made adaptive adjustments, such as the aforementioned correction. If either criterion is met, we consider the model to have successfully identified the issue as unreasonable.
The results are shown in Figure 4. It can be seen that even strong models like GPT-4 and Qwen-Max cannot fully identify the irrationality of the problem when directly facing the math problem. This indicates that this phenomenon is universal in large models.
Figure 4: The accuracy on UMP benchmark of different LLMs.
4 Do LLMs have the potential to discover and fix irrationality hidden in the problems?
Overview In this section, we illustrate the inherent potential of models to discern the irrational elements within a given problem. As evidenced by the experimental results in section3.2.1, models exhibit a latent capability to identify errors when directly confronted with them, suggesting that the requisite ability is present but requires activation. Our findings point to a promising low-cost intervention that markedly enhances model performance in addressing and processing unreasonable queries.
4.1 Stimulate critical thinking ability in the model
We believe that models may lack some criticality towards user-provided questions. When presented with mathematical problems, models typically proceed under the assumption that the user’s question itself is reasonable. This assumption leads models to engage in step-by-step problem-solving, culminating in the errors depicted in Figure 2. To mitigate this, we introduce a preemptive measure by incorporating a cautionary note within the prompt, alerting the model to the potential unreasonableness of the question. Consequently, this adjustment redefines the model’s task, expanding its role from merely providing solutions to also evaluating the reasonableness of the question posed.
To enhance models’ ability and encourage a critical assessment of the problems presented, we integrate an additional directive into the prompt: "If you encounter any questions that appear unreasonable or impractical, please highlight these concerns in your response." This addition serves as a crucial reminder, prompting the model to not only solve but also evaluate the feasibility and reasonableness of the questions it faces.
4.2 Explore the "Implicit errors" and the "Explicit errors"
Our investigation into the model’s capacity to discern the reasonableness of questions uncover a notable distinction in its evaluative effectiveness. We observe that models struggle with direct judgments of question validity due to the nature of certain inaccuracies requiring computational discovery, which we have termed "implicit errors." Conversely, "explicit errors" are identifiable through mere examination of the question’s text without necessitating any calculations. For instance, Example2 illustrates an explicit error where the misuse of "g" as a unit of length rather than weight in "140g long" clearly denotes an irrational statement. On the other hand, Example3 represents an implicit error, where the inconsistency in "tire needs to be replaced" mileage only becomes apparent through calculation, not through initial reading. This distinction underscores the complexity of automatically evaluating question reasonableness and highlights the need for nuanced model training strategies to address both explicit and implicit errors effectively.
Example2:A question with Explicit Error Shania is designing her own dress, and decides to make it a longer dress by extending the dress by 50% of its original length. She also adds 20g to the bottom of the dress with a lace trim. If the final design is 140g long then how long, in grams, was the dress in its original design?
Example3:A question with Implicit Error Every 2 miles a car drives, the tires rotate 725 times. Jeremy drives 400 miles a month. If each tire can undergo 10,440,000 rotations, how many years before the tire needs to be replaced if Jeremy needs to replace his tires after driving 300,000 miles?
Basic CoT Direct CCC(Ours)
Model Accuracy F1-score Accuracy F1-score Accuracy F1-score
ChatGLM3-6B 21 33.6 26 39.7 38 51.0
Llama2-7B 17 28.3 18 29.3 30 43.5
Baichuan2-13B 15 25.2 15 25.0 36 49.7
Qwen-Max 75 84.3 78 85.2 96 92.8
GPT-3.5 68 78.6 71 80.2 94 92.2
GPT-4 78 85.2 80 84.7 96 89.3
Table 2: Model performance comparison across different methods, which shows that our method achieves better results on models of all sizes. Accuracy is the performance on the 100 question unreasonable benchmark, f1-score is the score on 100 unreasonable questions and their corresponding 100 original reasonable questions. This score ensures that our template tries not to cause the model to judge reasonable questions as unreasonable.
4.3 A general approach to different types of problems
In response to the challenges identified in distinguishing between explicit and implicit errors, we propose a novel methodological framework: "Critical + Calculation + Conclusion." This framework initiates with the model engaging in a critical assessment of the posed question, followed by proceeding with the calculation as per standard practice. When an error is encountered during the calculation phase, the process is halted, and a conclusion is formulated, detailing the erroneous aspect of the question. This strategy offers dual benefits: for accurately formulated questions, the model efficiently generates correct answers; for unreasonable questions, it promptly identifies and articulates the specific issues. Thus, this approach significantly enhances the model’s capability to discern and address both correct and incorrect queries with improved precision and reliability.
Initially, we prompt the model with the instruction, "Please solve these problems with a critical thinking." This is immediately followed by, "Your response should commence with a step-by-step calculation process." Subsequently, the model is tasked with assessing the reasonableness of the problem, leading to a conclusion based on this evaluation. This critical juncture is articulated through the guideline, "If the problem is reasonable, furnish the correct answer. Should the problem exhibit inconsistencies, errors, elucidate these issues with clarity in your response." Finally, the model is instructed to "provide a succinct explanation," aiming to foster a deeper understanding of the problem-solving process and its outcomes.
4.4 Preference Improvement on CCC prompt
We compare out approach with Chain of Thought (CoT) and direct query plausibility method, which strategy explicitly prompts the model to assess the reasonableness of the question. We prompts LLMs in a Zero-shot manner(Lampert et al., 2009), enhancing the model’s reasoning by prefacing responses with "Let’s think step by step."(Kojima et al., 2022) . As illustrated in table 2, our Comparative Cognition Calibration (CCC) method outperforms both CoT and direct query methods across various model sizes, demonstrating superior results. The F1-score indicates that our model effectively addresses reasonable questions, showcasing its proficiency in discerning reasonable inquiries rather than indiscriminately identifying issues.
5 Conclusion
In this research, we make contributions to the field of large language models (LLMs) and their application in mathematical problem-solving. Firstly, we construct a novel dataset, the Unreasonable Math Problems (UMP) benchmark, which exposes a critical limitation of LLMs in detecting and addressing problems with inherent unreasonableness. This discovery underscores the models’ inability to perceive certain types of errors without explicit guidance. Secondly, we demonstrate that with appropriate prompting, LLMs can indeed identify and correct unreasonable elements within math problems. We design an effective prompt strategy that guides the models to not only solve rational problems but also critically evaluate and amend unreasonable ones in specific scenarios. This solution is a low-cost intervention that significantly enhances the models’ ability to handle unreasonable mathematical queries and reduce the hallucinatory content.
Appendix A Prompts for generating UMP
In this section, we outline the procedure employed to utilize GPT-4 for the creation of the Unreasonable Math Problems (UMP) benchmark. Provided below is the template prompt we used. To generate specific instances of the benchmark, one must simply choose an appropriate class from the options listed and insert it into the . The refined process is straightforward and systematic, enabling the efficient production of varied problem sets.

### v2

@@ Abstract
Large language models (LLMs) demonstrate substantial capabilities in solving math problems. However, they tend to produce hallucinations when given questions containing unreasonable errors. In this paper, we study the behavior of LLMs when faced with unreasonable math problems and further explore their potential to address these problems. We construct the Unreasonable Math Problem (UMP) benchmark to examine the error detection ability of LLMs. Experiments show that LLMs are able to detect unreasonable errors, but still fail in generating non-hallucinatory content. In order to improve their ability of error detection and correction, we further design a strategic prompt template called Critical Calculation and Conclusion(CCC). With CCC, LLMs can better self-evaluate and detect unreasonable errors in math questions, making them more reliable and safe in practical application scenarios.
Large Language Models Are Unconscious of Unreasonability
in Math Problems
Jingyuan Ma1 , Damai Dai1 , Lei Sha2, Zhifang Sui1 1National Key Laboratory for Multimedia Information Processing, School of Computer Science, Peking University 2Institute of Artificial Intelligence, Beihang University {majingyuan@bit.edu.cn}
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Ope, 2023; Gem, 2023) have significantly advanced the progress of artificial intelligence. These models demonstrate impressive performance across a wide range of downstream tasks, including solving math problems like GSM8K (Cobbe et al., 2021) and MATH(Hendrycks et al., 2021). In scenarios such as intelligent education, an LLM with strong math ability can be a powerful and versatile tool.
@@
Figure 1: Example inputs and outputs with a) direct mathematical problem prompts and b) prompts that signal the potential unreasonableness of the math problems. We can find that a) presents 12.5 trees but b) points out that this is not reasonable
@@
Considering the strong error detection capability in LLMs(Tyen et al., 2024), we further design a prompt template called Critical Calculation and Conclusion to stimulate and leverage the self-evaluation and critical thinking abilities in LLMs. To be specific, for each question, we urge the model to scrutinize the reasonability of each question. Then, it will directly answer reasonable questions, but correct and reevaluate unreasonable ones. With CCC, LLMs enhance their ability to identify and highlight unreasonable part within a problem. Simultaneously, this approach ensures the model’s performance remains unaffected when dealing with problems that are deemed reasonable.
Our contributions are summarized as follows: (1) We find that many LLMs’ performance significantly diminishes when encountering unreasonable problems, posing potential security risks. (2) We propose a semi-automatic data generation method and use it to develop the Unreasonable Math Problem (UMP) benchmark, which systematically assess the model’s ability in addressing unreasonable problems. (3) We introduce the CCC prompt template designed to stimulate the model’s self-detection capabilities. This method aids the model in identifying and rectifying unreasonable problems efficiently.
@@ 2 Explore how LLMs respond when faced with unreasonable problems
Overview Although the existing technology can achieve good results in solving mathematical reasoning problems, models are still prone to generate some inaccurate and confusing responses when confronted with nonsensical or unreasonable questions. As illustrated in Figure 1, to thoroughly evaluate the performance of large language models in the face of the above phenomenon, we construct a benchmark named Unreasonable Math Problems, containing unreasonable questions and test it on different models.
@@ 2.1 The UMP Benchmark
Our benchmark comprises 150 mathematically unreasonable problems and 150 original reasonable problems. Each original problem has been derived from a corresponding valid question sourced from the established GSM8K dataset (Cobbe et al., 2021), ensuring that the foundational structure of the problems remains stable. This subsection details the benchmark construction process.
@@ 2.1.1 Data Category
We identify and categorize five prevalent types of unreasonable math problems that commonly challenge large language models. These categories are: Undefined Variables, Illogical Scenarios, Incorrect Assumptions, Misinterpretation of Units, and Inconsistent Conditions. [^1] Due to space constraints, detailed descriptions of each category have been moved to the appendix A.
Figure 2: Our semi-automatic framework for data generation.
2.1.2 Data Generation
We adopt a methodology akin to that of MetaMath(Yu et al., 2023b), utilizing ChatGPT to facilitate the generation of data. The overall framework can be seen in Figure 2. By selecting a subset of problems from the GSM8K(Cobbe et al., 2021) dataset and carefully craft a rule set by different categories. We prompt the GPT4 to generate unreasonable problems and explanations based on the original ones. Next, we encode the input sentences (original and newly generated questions, denoted as $x1$ and $x2$ respectively) using a pre-trained language model such as BERT(Devlin et al., 2019), represented as $(x1,x2)$. We then select the $k-th$ largest value of the similarity function $f$, denoted as $f_{k}$. Here, $k$ is chosen as the minimum of the sentence number of $x_{1}$ and $x_{2}$, and we compute the cosine similarity as our similarity metric. If $f_{k}$ exceeds a predefined threshold $t$, we proceed to the next step; otherwise, we discard the generated question. In the process of generating the UMP benchmark, we chose to set the threshold $t$ to 0.75. However, for the category of Illogical Scenarios, we relaxed the threshold to 0.65 due to the content of the problem needs to be modified more.
2.2 Performance on different LLMs
2.2.1 Explore whether LLMs have the ability to judge unreasonablility
To evaluate the model’s ability to handle unreasonable problems, we first need to ensure its ability to identify unreasonable expressions. If the model cannot directly identify inaccuracies, its effectiveness in detecting errors within questions is compromised. Our initial analysis shows that errors like Incorrect Assumptions and Misinterpretation of Units often arise from common sense misunderstandings, such as the impossibility of negative human attributes or recognizing a gram as a unit of weight. We created a small benchmark with 30 items in each category to test the model’s ability to recognize such irrational statements. The methodology and results are detailed in the Appendix D. The results in Table 1 show that models are generally capable of assessing the rationality of statements.
Model/Accuracy IA MU
ChatGLM3-6B 80.0 76.6
Llama2-7B-chat 80.0 73.3
Baichuan2-13B-chat 83.3 83.3
Qwen-Max 100.0 96.6
GPT-3.5 93.3 100
GPT-4 100.0 96.6
Table 1: The accuracy with which various models successfully identified irrationality when directly confronted with errors categorized as Incorrect Assumptions (IA) and Misinterpretation of Units (MU). Specific definitions of the categories can be found in the Appendix A
2.2.2 Performance on UMP benchmark
We adopt the common Chain of Thought (CoT) method as a prompt and set temperature to 0.7. To evaluate the model’s output, we first look for keywords or expressions like “unreasonable”. Next, we assess whether the model has made adaptive adjustments, such as the aforementioned correction. If either criterion is met, we consider the model to have successfully identified the issue as unreasonable.
The results are shown in Figure 3. To ensure that any shortcomings in identifying illogical aspects were not simply due to a lack of mathematical reasoning abilities, we conducted an evaluation of the selected models’ accuracy in solving the original problems. As depicted by the green line in Figure 3, there is a discernible discrepancy between the models’ performance in solving problems correctly(facing the original problems) and their success in identifying unreasonable elements within unreasonable problems, which is shown in blue. This discrepancy persists even when accounting for the potential limitations in mathematical reasoning. We assume that this gap, evident between the green and blue lines in the figure, could potentially be narrowed by leveraging and enhancing the models’ existing capabilities.
The performance indicated by the yellow line reflects the models’ proficiency in deeming a problem reasonable when presented with original problems. This suggests that the models will not find fault in the face of such reasonable problems.
Figure 3: The Performance of different LLMs. The gap between the green and blue lines is the capability of the model we manage to stimulate.
3 Do LLMs have the potential to discover irrationality hidden in the problems?
Overview In this section, we illustrate the inherent potential of models to discern the irrational elements within a given problem. As evidenced by the experimental results in section2.2.1, models exhibit a latent capability to identify errors when directly confronted with them, only need to stimulate it through specific means.
3.1 Stimulate Critical Thinking Ability in the Model
We believe that models may lack some criticality towards user-provided questions. We propose a novel methodological framework: “Critical + Calculation + Conclusion.” This framework initiates with the model engaging in a critical assessment of the posed question, followed by proceeding with the calculation as per standard practice. When an error is encountered during the calculation phase, the process is halted, and a conclusion is formulated, detailing the erroneous aspect of the question. This strategy offers dual benefits: for accurately formulated questions, the model efficiently generates correct answers; for unreasonable questions, it promptly identifies and articulates the specific issues. The template for our CCC can be found in the appendix E.
Basic CoT Direct CCC(Ours)
Model Acc. F1 Acc. F1 Acc. F1
ChatGLM3-6B 20.6 33.2 26.3 39.7 38.6 53.2
Llama2-7B 14.0 23.2 18.3 29.3 27.3 34.3
Baichuan2-13B 14.6 24.3 18.6 29.5 35.3 49.3
Qwen-Max 70.6 81.2 78.0 85.2 94.0 92.8
GPT-3.5 66.0 78.0 69.3 79.4 89.3 90.5
GPT-4 74.6 83.3 80.0 84.7 94.6 92.2
Table 2: Model performance comparison across different methods. Accuracy is the performance on the unreasonable problems, F1-score is the score on both unreasonable questions and their corresponding original reasonable questions.
3.2 Preference Improvement on CCC prompt
We compare out approach with Chain of Thought (CoT) and direct query plausibility method, which strategy explicitly prompts the model to assess the reasonableness of the question. We prompts LLMs in a Zero-shot manner(Kojima et al., 2022). As illustrated in Table 2, our critical calculation and conclusion (CCC) method outperforms both CoT and direct query methods across various model sizes, demonstrating superior results. The F1-score indicates that our model effectively addresses reasonable questions, showcasing its proficiency in discerning reasonable inquiries rather than indiscriminately identifying issues.
4 Conclusion
In this research, we make contributions to the field of LLMs’ application in mathematical problem-solving. Firstly, we propose a semi-automatic data generation method and use it to construct Unreasonable Math Problems (UMP) benchmark, which exposes a critical limitation of LLMs in detecting and addressing problems with inherent unreasonableness. This discovery underscores the models’ inability to perceive certain types of errors without explicit guidance. Secondly, we demonstrate that with appropriate prompting, LLMs can indeed identify and correct unreasonable elements within math problems. We design an effective prompt strategy that guides the models to not only solve rational problems but also critically evaluate and amend unreasonable ones in specific scenarios. This solution is a low-cost intervention that significantly enhances the models’ ability to handle unreasonable mathematical queries and reduce the hallucinatory content.
Limitations
Our work has several limitations. Firstly, due to funding constraints, our scale is relatively small in order to ensure the data quality of our benchmark. However, we elaborate on the semi-automatic data synthesis method in the paper, which shows that with sufficient funding and human assistance, our method can easily scale up and generate more high-quality data. Secondly, our analysis primarily compares the traditional method of solving mathematical problems and does not definitively establish whether our method is orthogonal to other problem-solving approaches. Future research can test more models and methods based on our benchmark to explore this further.
Ethical Considerations
Our paper explores how llms perform in the face of unreasonable mathematical problems. Sometimes the model may not realize the irrationality in the math problem, so it may produce wrong answers and mislead the user.
Appendix A Category of UMP Benchmark
Below is the description of the five categories. Each of our data pieces is made up of five parts. An original question $q$ paired with its corresponding answer $a$, along with a designated type of unreasonableness $t$ we aim to investigate, we systematically construct a rephrased, unreasonable question $q^{\prime}$ and provide an accompanying explanation $e$. Consequently, our benchmark comprises quintuples of the form $(q,a,q^{\prime},t,e)$.
@@ Appendix A Category of UMP Benchmark
Appendix B Semi-automatic framework for generating UMP
B.1 Prompt for Semi-automatic framework
In this section, we outline the procedure employed to utilize GPT-4 for the creation of the Unreasonable Math Problems (UMP) benchmark. Firstly, we provide the model with detailed rule descriptions, instructing it on how to transform original questions into unreasonable ones, category by category. Then, we supply the model with the original questions, allowing it to autonomously choose one of the several invalid categories for rewriting. We do not specify the category because not all questions are suitable for transformation into certain categories. For example, errors such as “Misinterpretation of Units” require the original question to contain content with units, so not all questions can be converted into this type of invalidity.